OpenStudy (anonymous):

HELPPPP PLEASE! EASY MEDAL. Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 a) -4x - 2y = 10 -4y = 4 B) -4x - 5y = -1 -7y = 5 C) 3x + 2y = 6 7x = 12 D)4x + 2y = 10 8x = 16

OpenStudy (wiggler):

Starting with the original equations, taking 2 of the second eqn away from the first: \[4x-2y-2(2x+y)=6-2(5)\] \[-4y=-4\] \[y=1\] And putting that into the second equation, x=1

OpenStudy (wiggler):

Sorry, x=2

OpenStudy (wiggler):

So putting these values for x and y into the given equations, we find that D is the only one that works.